Given the following information, what is the mass attenuation coefficient of this material? What is material?

Mean density = 2.33 g/cm3

Io = 2 Mev

Thickness (cm)

Measured gamma emitted (Mev)

2.0

1.65

2.3

1.53

3.0

1.40

3.1

1.42

4.2

1.29

4.2

1.31

5.0

1.16
Answer #1

Using equation,

ln(I) = ln(Io) - ux

with,

I = inensity data measured

Io = incident intensity

u = linear attenuation coefficient

x = thickness

plot, ln(I) on y-axis and x on x-axis

slope of curve = -u = -0.0518

Thus, mass attenuation coefficient = 0.0518/2.33 = 0.022 cm^2/g

plot In(l) vs x 0.6 0.5 0.4 E 0.3 0.2 0.1 yー-0.0518x+0.5337 2 2.3 3 3.1 4.2 4.2 5 x (cm)